OpenStudy (anonymous):

find all the trigonometric ratios for (DELTA) = 5pi/3

OpenStudy (anonymous):

locate \[\frac{5\pi}{3}\] on the unit circle on last page of attached cheat sheet.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the first coordinate is cosine the second coordinate is sine sine over cosine is tangent etc

OpenStudy (anonymous):

you should see the coordinate is \[(\frac{1}{2},-\frac{\sqrt{3}}{2})\] so \[\cos(\frac{5\pi}{3})=\frac{1}{2}\] \[\sin(\frac{5\pi}{3})=-\frac{\sqrt{3}}{2}\] tec

OpenStudy (anonymous):

*etc

OpenStudy (anonymous):

ok i think i understand that...what would i do to find all the trig values for delta = 5pi/3 ..could you explain that...im still confused

OpenStudy (anonymous):

hey...i figured it out ha thank you!